A 1.5 L solution of chemotherapeutic drug with a molecular mass of 539 and a concentration of 0.05 M is being administered to a cancer patient. How many grams of the drug enter the patients body?
1
Expert's answer
2015-09-01T03:06:08-0400
Find the number of moles of the drug: n = c*V = 0.05M*1.5L = 0.075mol.
The number of grams of the drug will be: m = 0.075mol*539 = 40.425g